# CLIP-HandID: Vision-Language Model for Hand-Based Person Identification

### v1

@@ Abstract
This paper introduces a new approach to person identification based on hand images, designed specifically for criminal investigations. The method is particularly valuable in serious crimes like sexual abuse, where hand images are often the sole identifiable evidence available. Our proposed method, CLIP-HandID, leverages pre-trained foundational vision-language model, particularly CLIP, to efficiently learn discriminative deep feature representations from hand images given as input to the image encoder of CLIP using textual prompts as semantic guidance. We propose to learn pseudo-tokens that represent specific visual contexts or appearance attributes using textual inversion network since labels of hand images are indexes instead text descriptions. The learned pseudo-tokens are incorporated into textual prompts which are given as input to the text encoder of the CLIP to leverage its multi-modal reasoning to enhance its generalization for identification. Through extensive evaluations on two large, publicly available hand datasets with multi-ethnic representation, we show that our method substantially surpasses existing approaches. The source code is available at https://github.com/nathanlem1/CLIP-HandID.
@@ I Introduction
Biometric identification, which verifies individuals based on physical or behavioral traits, has gained significant interest for various applications. Among primary biometric traits, hand images provide distinctive features for person identification [1, 2, 3, 4, 5]. Unlike other biometric modalities, hand images exhibit lower variability while maintaining robust and diverse characteristics that remain stable in adulthood [2, 3, 6, 7]. This stability makes hand images a promising candidate for biometric analysis, particularly in forensic investigations where uncontrolled environments are common. In serious crimes like sexual abuse, hand images are often the sole identifiable evidence available, underscoring their investigative value.
Previous studies [8, 6] have explored hybrid approaches combining traditional machine learning and deep learning techniques for hand-based person identification. In [8], researchers trained a convolutional neural network (CNN) on RGB hand images and subsequently used the network as a feature extractor. The extracted CNN features were then classified using support vector machine (SVM) classifiers. Similarly, [6] adopted a comparable methodology but incorporated an additional data modality by fusing near-infrared (NIR) images with RGB data. However, these approaches lack a true end-to-end learning framework. End-to-end approaches have also been introduced in [2, 3, 9]; however, they are based on uni-modal architectures, processing only images.
@@
    We learn pseudo-tokens that represent specific visual contexts using textual inversion network since labels of hand images are indexes instead of text descriptions.
@@ II Proposed Method
While CLIP operates effectively in standard classification by reformulating labels as text prompts (yielding aligned visual $I_{i}$ and text $T_{i}$ embeddings), its direct application to hand-based person identification is problematic due to the non-semantic nature of indexes as hand labels instead of text descriptions, which serve as mere instance identifiers. Therefore, we design textual inversion network to directly invert an image (visual embedding $I_{i}$) into a unique pseudo-token $S_{*}$, which can be incorporated into a textual prompt as ”A photo of a $S_{*}$ hand” and then given as input to the text encoder of the CLIP, as shown in Fig. 1. It is important to highlight that this pseudo-token does not correspond to any real word; instead, it serves as an abstract representation within the token embedding space. The learned pseudo-token effectively encode both global visual patterns and fine-grained visual characteristics of the original image. Furthermore, the learned pseudo-tokens by the textual inversion network are personalized textual descriptions for individual subjects which enable or guide the model to focus on semantically relevant visual regions (hand structure) that align precisely with the prompt context. We use a lightweight model for the inversion network consisting of three fully-connected (FC) layers with GELU activation functions, layer normalization and dropout with a probability of 0.5 after each layer except the last layer. Only the layer normalization is used just after the last layer.
The visual embedding $I_{i}$ is given as input to a classification layer which is implemented using a FC layer followed by a softmax function. We also include a batch normalization (BN) to reduce possible over-fitting, as shown in Fig. 1. The classification layer predicts the identity (ID) of each input image.
@@
where $I_{i}=\theta_{I}(\mathbf{x}_{i})$ and $T_{i}=\theta_{T}(\mathbf{t}_{i})$ are the image feature and the text feature, respectively, projected into a cross-modal embedding space, $\tau$ represents the temperature (we set $\tau=1$), and $\text{sim}(u,v)=\frac{u\cdot v}{\|u\|_{2}\|v\|_{2}}$ is the cosine similarity. Then, the image-to-text contrastive loss $\mathcal{L}_{i2t}$ is calculated as:
@@
where the numerators in (4) and (5) represent the similarity scores between matched embedding pairs, while the denominators aggregate all similarity measures relative to the anchor $I_{i}$ or $T_{i}$.
@@ IV Conclusion
In this work, we propose a novel hand-based person identification framework tailored for forensic applications, where hand imagery often serves as the primary evidence in severe criminal cases such as sexual assault. Our proposed method, CLIP-HandID, innovatively adapts the CLIP vision-language model to this domain by developing a mechanism to bridge the gap between numeric identity labels and semantic visual representations. The core innovation involves learning task-specific pseudo-tokens through textual inversion, which encode distinctive hand attributes (e.g., vein patterns, knuckle geometry, etc.) into adaptable text prompts. These learned pseudo-tokens are integrated into structured semantic prompts (e.g., ”A photo of a $S_{*}$ hand”, where $S_{*}$ is the learnt pseudo-token) that guide CLIP’s cross-modal alignment. The image encoder processes input hand images while the text encoder interprets our enhanced prompts, enabling more discriminative feature learning through the model’s pre-trained multi-modal reasoning capabilities. Comprehensive experiments on two large-scale, ethnically diverse hand datasets demonstrate that CLIP-HandID significantly outperforms current state-of-the-art methods in recognition accuracy. This advancement is particularly crucial for forensic investigations where conventional biometric features may be unavailable.

### v2

@@ Abstract
This paper introduces a novel approach to person identification using hand images, designed specifically for criminal investigations. The method is particularly valuable in serious crimes such as sexual abuse, where hand images are often the only identifiable evidence available. Our proposed method, CLIP-HandID, leverages a pre-trained foundational vision-language model - CLIP - to efficiently learn discriminative deep feature representations from hand images (input to CLIP’s image encoder) using textual prompts as semantic guidance. Since hand images are labeled with indexes rather than text descriptions, we employ a textual inversion network to learn pseudo-tokens that encode specific visual contexts or appearance attributes. These learned pseudo-tokens are then incorporated into textual prompts, which are fed into CLIP’s text encoder to leverage its multi-modal reasoning and enhance generalization for identification. Through extensive evaluations on two large, publicly available hand datasets with multi-ethnic representation, we demonstrate that our method significantly outperforms existing approaches. The source code is available at https://github.com/nathanlem1/CLIP-HandID.
@@ I Introduction
Biometric identification, which verifies individuals based on physical or behavioral traits, has gained significant interest for various applications. Among primary biometric traits, hand images provide distinctive features for person identification [1, 2, 3, 4, 5]. Unlike other biometric modalities, hand images exhibit lower variability while maintaining robust and diverse characteristics that remain stable in adulthood [2, 3, 6, 7]. This stability makes hand images a promising candidate for biometric analysis, particularly in forensic investigations where uncontrolled environments are common. In serious crimes like sexual abuse, hand images are often the only identifiable evidence available, underscoring their investigative value.
Previous studies [8, 6] have explored hybrid approaches combining traditional machine learning and deep learning techniques for hand-based person identification. In [8], researchers trained a convolutional neural network (CNN) on RGB hand images and subsequently used the network as a feature extractor. The extracted CNN features were then classified using support vector machine (SVM) classifiers. Similarly, [6] adopted a comparable methodology but incorporated an additional data modality by fusing near-infrared (NIR) images with RGB data. However, these approaches lack a true end-to-end learning framework. End-to-end approaches have also been introduced in [2, 3, 9]; however, they are based on uni-modal architectures and process only images.
@@
    We learn pseudo-tokens that encode specific visual contexts using textual inversion network since hand images are labeled with indexes rather than text descriptions.
@@ II Proposed Method
While CLIP operates effectively in standard classification by reformulating labels as text prompts (yielding aligned visual $I_{i}$ and text $T_{i}$ embeddings), its direct application to hand-based person identification is problematic due to the non-semantic nature of indexes as hand image labels instead of text descriptions, which serve as mere instance identifiers. Therefore, we design textual inversion network to directly invert an image (visual embedding $I_{i}$) into a unique pseudo-token $S_{*}$, which can be incorporated into a textual prompt as ”A photo of a $S_{*}$ hand” and then given as input to the text encoder of the CLIP, as shown in Fig. 1. It is important to highlight that this pseudo-token does not correspond to any real word; instead, it serves as an abstract representation within the token embedding space. The learned pseudo-token effectively encode both global visual patterns and fine-grained visual characteristics of the original image. Furthermore, the learned pseudo-tokens by the textual inversion network are personalized textual descriptions for individual subjects which enable or guide the model to focus on semantically relevant visual regions (hand structure) that align precisely with the prompt context. We use a lightweight model for the inversion network consisting of three fully-connected (FC) layers with GELU activation function, layer normalization and dropout with a probability of 0.5 after each layer except the last layer. Only the layer normalization is used just after the last layer.
The visual embedding $I_{i}$ is given as input to a classification layer, which is implemented using a FC layer followed by a softmax function. We also include a batch normalization (BN) to reduce possible over-fitting, as shown in Fig. 1. The classification layer predicts the identity (ID) of each input image.
@@
where $I_{i}=\theta_{I}(\mathbf{x}_{i})$ and $T_{i}=\theta_{T}(\mathbf{t}_{i})$ are the image feature and the text feature, respectively, projected into a cross-modal embedding space, and $\text{sim}(u,v)=\frac{u\cdot v}{\|u\|_{2}\|v\|_{2}}$ is the cosine similarity. Then, the image-to-text contrastive loss $\mathcal{L}_{i2t}$ is calculated as:
@@
where the numerators in (4) and (5) represent the similarity scores between matched embedding pairs, while the denominators aggregate all similarity measures relative to the anchor $I_{i}$ or $T_{i}$, and $\tau$ represents the temperature, which we set $\tau=1$.
@@ IV Conclusion
In this work, we propose a new hand-based person identification framework tailored for forensic applications, where hand imagery often serves as the primary evidence in severe criminal cases such as sexual assault. Our proposed method, CLIP-HandID, innovatively adapts the CLIP vision-language model to this domain by developing a mechanism to bridge the gap between numeric identity labels and semantic visual representations. The core innovation involves learning task-specific pseudo-tokens through textual inversion, which encode distinctive hand attributes (e.g., vein patterns, knuckle geometry, etc.) into adaptable text prompts. These learned pseudo-tokens are integrated into structured semantic prompts (e.g., ”A photo of a $S_{*}$ hand”, where $S_{*}$ is the learnt pseudo-token) that guide CLIP’s cross-modal alignment. The image encoder processes input hand images while the text encoder interprets our enhanced prompts, enabling more discriminative feature learning through the model’s pre-trained multi-modal reasoning capabilities. Comprehensive experiments on two large-scale, ethnically diverse hand datasets demonstrate that CLIP-HandID significantly outperforms current state-of-the-art methods in recognition accuracy. This advancement is particularly crucial for forensic investigations where conventional biometric features may be unavailable.